Given cos 0=-4/5, and the angle 0 is in quadrant two, find the value for sin 0.
Answer by lwsshak3(11628)

Given cos 0=-4/5, and the angle 0 is in quadrant two, find the value for sin 0
***
Given angle O is from a 3-4-5 reference right triangle in quadrant II where cos<0 and sin>0.
If cos O=-4/5, sin O=3/5